Simultaneous equations - one linear, one quadratic
x^2 + y^2 = 25
x - y + 1 = 0

Answer by Alan3354(69443)

Simultaneous equations - one linear, one quadratic
x^2 + y^2 = 25
x - y + 1 = 0 --> y = x+1
Sub for y in the 1st equation.
Solve for x
etc